OpenStudy (anonymous):

Question for fun, one sec while i type it up. Question for fun, one sec while i type it up. @Mathematics

OpenStudy (anonymous):

Let K be the set of all positive integers whose prime factorization only has 2's, 3's and 5's. Evaluate:\[\sum_{x\in K}\frac{1}{x}=\frac{1}{2}+\frac{1}{3}+\frac{1}{4}+\frac{1}{5}+\frac{1}{6}+\frac{1}{8}+\frac{1}{9}+\frac{1}{10}+\frac{1}{12}+\cdots\]

OpenStudy (anonymous):

Solution.